# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.1 Proofs of the Auxiliary Lemmas and Resutls
$\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{F}=\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{F}\|\bm{\Gamma}^{-1}_{1,p-1}\|^{2}_{\mathrm{op}}+\|\bm{\Gamma}^{-1}_{1,p-1}-\bm{\Gamma}^{-1}_{2,p-1}\|^{2}_{F}\|\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}$.
There are a total of $T-p$ blocks of $\bm{\xi}_{1,p}$ in $\bm{\Upsilon}_{1,T}$. Thus We have $(T-p)\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}\implies\frac{T-p}{T}\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{F}\leq\frac{1}{T}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$. Furthermore, $\|\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\xi}_{1,p-1}\|^{2}_{\mathrm{op}}+\|\bm{\xi}_{1,p-1}-\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}$.
In $\bm{\Upsilon}_{1,T}$, there are $T/p$ diagonal blocks of $\bm{\Upsilon}_{1,p-1}$’s assuming $p$ divides $T$, and thus $(T/p)\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$. Hence, $\frac{T/p}{T}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}\leq\frac{1}{T}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$.
For $p$ not dividing $T$ is redundant in an asymptotic regime. ∎
@@
We assume $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|_{\mathrm{F}},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|_{\mathrm{F}},\|\bm{U}_{1,j}-\bm{U}_{2,j}\|_{\mathrm{F}},\|\bm{V}_{1,j}-\bm{V}_{2,j}\|_{\mathrm{F}}\}\leq\epsilon$. Since, $\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{j-1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j}\|^{2}_{\mathrm{op}}$ and $\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{D}_{j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{p}^{-1}\|_{\mathrm{op}}$, we have $\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{L}_{\ell,j}\|^{2}_{\mathrm{op}}\leq M_{\ell,U}$, $\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{\ell,p}^{-1}\|_{\mathrm{op}}=M_{\ell,V}$ and $(1+\|\bm{C}^{-1}_{1,p}\|_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{L}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{p}^{-1}\|_{\mathrm{op}})^{p}\leq M_{\ell,C,V,U}$, we have,
$\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}$
@@
$\qquad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\{p^{4}M_{1,U}M_{1,V}M_{1,C,U,V}+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}$
$\qquad+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{F}M_{2,U}M_{1,V}M_{1,C,U,V}\}]$
And,
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{F}$
@@
where $M_{P,1}=[(p+1)+p^{3}\{M_{1,U}M_{1,V}M_{2,U}M_{2,V}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(M_{1,V}+M_{2,U})\}+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\{p^{4}M_{1,U}M_{1,V}M_{1,C,U,V}+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{F}M_{2,U}M_{1,V}M_{1,C,U,V}\}]$ and $M_{P,2}=[2p^{2}(M_{1,U}M_{1,V}M_{1,C,U,V}+p)+2p\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}+2p\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{F}M_{2,U}M_{1,V}M_{1,C,U,V}]$.
While bounding around the truth, we let the eigenvalues of $\bm{\Omega}_{1}$ be bounded between two fixed constants. Also let $\|\bm{L}_{1,j}\|^{2}_{\mathrm{op}},\|\bm{K}_{1,j}\|^{2}_{\mathrm{op}}$ be bounded by again a fixed constant. This puts an upper bound on $\bm{C}_{p}^{-1}$ as well. We have $\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\lesssim\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\}$.
We can similarly show that, $\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\lesssim\{\|\bm{D}_{j-1,1}-\bm{D}_{j-1,2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}$
We have, using the recurrence relations, that $\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{\mathrm{op}}\leq\|\bm{D}_{1,j-2}-\bm{D}_{2,j-2}\|^{2}_{\mathrm{op}}+\|\bm{U}_{1,j-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j-1}\|^{2}_{\mathrm{op}}\|\|\bm{V}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-2}^{-1}-\bm{C}_{2,j-2}^{-1}\|^{2}_{\mathrm{op}}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,j-1}-\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j-1}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j-1}-\bm{V}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}})$.
Then, we have $\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}\leq\epsilon,\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}\leq\epsilon$ if $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}\lesssim\epsilon$ and $\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{op}}\lesssim\epsilon$. The constants will depend on the bounds on the truth.
@@
In the sieve, we have $\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\leq u_{T}\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\}$ for some $u_{T}$, a polynomial in $T$.
We can similarly show that, $\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\leq v_{T}\{\|\bm{D}_{j-1,1}-\bm{D}_{j-1,2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}$ for some $v_{T}$ which is polynomial in $T$.
Then, we have $\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}\leq\epsilon,\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}\leq\epsilon$ if $\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}\leq\min\left\{\frac{\epsilon}{u_{T}},\frac{\epsilon}{v^{\prime}_{T}}\right\},\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\leq\min\left\{\frac{\epsilon}{pu_{T}},\frac{\epsilon}{pv^{\prime}_{T}}\right\}$. In summary, we need $\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{op}}\leq o_{T}\epsilon$, for some polynomial in $T$, $o_{T}$ within ${\cal G}_{T}$ for the above to hold.
Finally, $\|\bm{D}_{1,0}-\bm{D}_{2,0}\|^{2}_{\mathrm{op}}\leq c_{T}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}$ for some $c_{T}$ which is polynomial in $T$ and the bound for $\|\bm{U}_{1,j-1}-\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}$ is established above.
Thus, $\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\leq v^{\prime}_{T}\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j-1}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\}$ for some $v^{\prime}_{T}$ which is polynomial in $T$
Within ${\cal G}_{T}$, if $a_{T,1},a_{T,2},b_{T},c_{T}$ are all polynomials in $T$, we have $M_{P,1}=c_{T}$ and $M_{P,2}=e_{T}$ which are also some polynomials in $T$.
Setting, $\epsilon=\delta_{T}/[T^{2}c_{T}+(T-p+1)a_{T,1}e_{T}]$, we have $\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|_{\mathrm{F}}\leq\delta_{T}$, where $\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}\leq\epsilon,\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}\leq\epsilon,\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{F}\leq\epsilon,\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\leq\epsilon$.
Then using  (A.4), we have the final result.

### v2

@@ A.1 Proofs of the Auxiliary Lemmas
Note that
$\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{\mathrm{F}}=\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{\mathrm{F}}\|\bm{\Gamma}^{-1}_{1,p-1}\|^{2}_{\mathrm{op}}+\|\bm{\Gamma}^{-1}_{1,p-1}-\bm{\Gamma}^{-1}_{2,p-1}\|^{2}_{\mathrm{F}}\|\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}.$
As there are $T-p$ blocks of $\bm{\xi}_{1,p}$ in $\bm{\Upsilon}_{1,T}$, $(T-p)\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{\mathrm{F}}\leq\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ which implies that $(1-p/T)\|\bm{\xi}_{1,p}-\bm{\xi}_{2,p}\|^{2}_{\mathrm{F}}\leq T^{-1}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$. Moreover, we have the bound $\|\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\xi}_{1,p-1}\|^{2}_{\mathrm{op}}+\|\bm{\xi}_{1,p-1}-\bm{\xi}_{2,p-1}\|^{2}_{\mathrm{op}}$.
We assume without loss of generality that $p$ divides $T$; the contribution from the remainder may be neglected since $p/T\to 0$. As there are $T/p$ diagonal blocks of $\bm{\xi}_{1,p}$ in $\bm{\Upsilon}_{1,p-1}$,
$p^{-1}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}\leq T^{-1}(T/p)\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}\leq T^{-1}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}},$
establishing the claim. ∎
@@
Let
$\max\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|_{\mathrm{F}},\|\bm{L}_{1,j}-\bm{L}_{2,j}\|_{\mathrm{F}},\|\bm{U}_{1,j}-\bm{U}_{2,j}\|_{\mathrm{F}},\|\bm{V}_{1,j}-\bm{V}_{2,j}\|_{\mathrm{F}}\}\leq\epsilon.$
Since, $\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{j-1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j}\|^{2}_{\mathrm{op}}$ and $\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{D}_{j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{p}^{-1}\|_{\mathrm{op}}$, we have
$\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}$ $\leq\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{L}_{\ell,j}\|^{2}_{\mathrm{op}}\leq M_{\ell,U},$
$\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}$ $\leq\|\bm{C}_{\ell,p}^{-1}\|_{\mathrm{op}}=M_{\ell,V},$
$(1+\|\bm{C}^{-1}_{1,p}\|_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{L}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{p}^{-1}\|_{\mathrm{op}})^{p}$ $\leq M_{\ell,C,V,U},$
we have,
$\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}$
@@
$\qquad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\{p^{4}M_{1,U}M_{1,V}M_{1,C,U,V}$
$\qquad+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{\mathrm{F}}M_{2,U}M_{1,V}M_{1,C,U,V}\}]$
and
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$
@@
where
$M_{P,1}$ $=[(p+1)+p^{3}\{M_{1,U}M_{1,V}M_{2,U}M_{2,V}$
$\quad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(M_{1,V}+M_{2,U})\}$
$\quad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\{p^{4}M_{1,U}M_{1,V}M_{1,C,U,V}$
$\quad+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}$
$\quad+2p^{2}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{\mathrm{F}}M_{2,U}M_{1,V}M_{1,C,U,V}\}]$
and
$M_{P,2}$ $=[2p^{2}(M_{1,U}M_{1,V}M_{1,C,U,V}+p)$
$\quad+2p\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}+2p\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{\mathrm{F}}M_{2,U}M_{1,V}M_{1,C,U,V}].$
While bounding around the truth, we let the eigenvalues of $\bm{\Omega}_{1}$ be bounded between two fixed constants. Also let $\|\bm{L}_{1,j}\|^{2}_{\mathrm{op}}$ and $\|\bm{K}_{1,j}\|^{2}_{\mathrm{op}}$ be also bounded by a fixed constant. This puts an upper bound on $\bm{C}_{p}^{-1}$ as well. Then we have
$\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\lesssim\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}.$
We can similarly show that
$\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\lesssim\|\bm{D}_{j-1,1}-\bm{D}_{j-1,2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}.$
The recurrence relations leads to
$\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{\mathrm{op}}$
$\leq\|\bm{D}_{1,j-2}-\bm{D}_{2,j-2}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{U}_{1,j-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j-1}\|^{2}_{\mathrm{op}}\|\|\bm{V}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-2}^{-1}-\bm{C}_{2,j-2}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,j-1}-\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j-1}\|^{2}_{\mathrm{op}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\|\bm{V}_{1,j-1}-\bm{V}_{2,j-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}).$
Then $\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}\lesssim\epsilon$ and $\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}\lesssim\epsilon$ if
$\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}\lesssim\epsilon,\;\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\lesssim\epsilon,\;\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{op}}\lesssim\epsilon.$
The implicit constants of proportionality will depend on the bounds for the truth.
@@
In the sieve, we have
$\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\leq u_{T}\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\}$
for some polynomial $u_{T}$ in $T$. We can similarly show that
$\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\leq v_{T}\{\|\bm{D}_{j-1,1}-\bm{D}_{j-1,2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}$
for some polynomial $v_{T}$ in $T$.
Then we have $\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}\leq\epsilon,\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}\leq\epsilon$ if
$\max\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}},\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}\}$ $\leq\min\big{\{}\frac{\epsilon}{u_{T}},\frac{\epsilon}{v^{\prime}_{T}}\big{\}},$
$\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}$ $\leq\min\big{\{}\frac{\epsilon}{pu_{T}},\frac{\epsilon}{pv^{\prime}_{T}}\big{\}}.$
The required condition will hold if $\|\bm{L}_{1,k}-\bm{L}_{2,k}\|^{2}_{\mathrm{op}}\leq o_{T}\epsilon$ for some polynomial $o_{T}\in{\cal G}_{T}$.
Finally, $\|\bm{D}_{1,0}-\bm{D}_{2,0}\|^{2}_{\mathrm{op}}\leq c_{T}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}$ for some $c_{T}$ which is polynomial in $T$ and the bound for $\|\bm{U}_{1,j-1}-\bm{U}_{2,j-1}\|^{2}_{\mathrm{op}}$ is established above. Thus
$\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\leq v^{\prime}_{T}\{\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}+\|\bm{K}_{1,j}-\bm{K}_{2,j}\|^{2}_{\mathrm{op}}+\sum_{k=1}^{j-1}\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\}$
for some $v^{\prime}_{T}$ which is polynomial in $T$
If $a_{T,1},a_{T,2},b_{T},c_{T}$ are all polynomials in $T$ belonging to ${\cal G}_{T}$, we have $M_{P,1}=c_{T}$ and $M_{P,2}=e_{T}$ also polynomials in $T$.
Setting, $\epsilon=\delta_{T}/[T^{2}c_{T}+(T-p+1)a_{T,1}e_{T}]$, we have $\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|_{\mathrm{F}}\leq\delta_{T}$, where
$\max(\|\bm{V}_{1,k}-\bm{V}_{2,k}\|_{\mathrm{F}},\|\bm{U}_{1,k}-\bm{U}_{2,k}\|_{\mathrm{F}},\|\bm{L}_{1,k}\bm{L}_{1,k}^{\mathrm{T}}-\bm{L}_{2,k}\bm{L}_{2,k}^{\mathrm{T}}\|_{\mathrm{F}},\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}})\leq\epsilon.$
Then, we reach the conclusion using (A.8). ∎